OpenStudy (anonymous):

What is the simplified form of x over x squared minus y squared + y over x squared minus y squared ? Help Please!?!

OpenStudy (anonymous):

is it this:\[\frac{ x }{ x^2-y^2 }+\frac{ y }{ x^2-y^2 }\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

they have the same denominator so you can write it as:\[\frac{ x+y }{ x^2-y^2 }\]

OpenStudy (anonymous):

if you factor the denominator you get:\[\frac{ (x+y) }{ (x+y)(x-y) }\]

OpenStudy (anonymous):

divide the top and bottom by x+y so it simplifies to:\[\frac{ 1 }{ x-y }\]

OpenStudy (anonymous):

Yes, I got the same thing. Thank you!